Student Instructions:

After studying this Module 11, answer the following questions in this Journal:

What is relative dating, and what can it tell us?
What is absolute dating, and what is it based on? What can absolute dating tell us?
What is the principle of uniformitarianism? Why does the principle of uniformitarianism lead to the conclusion that Earth's age is great?
What is the principle of superposition, and what is its importance? Who is credited with discovering it?
What do unconformities represent?
Why is correlation needed to construct a complete geologic history of a region? How does correlation work, e.g., in the Colorado Plateau region?
If you start with an element with 1,000,000 parent atoms, how many parents and how many daughters will there be after 6 half-lives?
What is Carbon-14 dating useful for in terms of age and type of sample?
Answer #1

1)Relative dating indicates the age of formation whether younger or older with the surrounding formation  

2)Absolute dating tells about actual age of formation when it was formed .it is based on radiiometric dating.It can tell us actual age ,when it was formed and how it was formed.

3)Principles of uniformtarnism tells us the processes that occurring today that has been happening similarly in ancient past.The conclusions derived from principle of uniformtarnism is present is key to past lead to Earth 's age is great.

4)Principle of superposition tells in a undisturbed sequence the younger age sediments deposit above older age sediments. This was discovered by scientists Steno.

5)Unconformity represents is area of erosion and non deposition.

6)Principle of superposition, principle of cross cutting relationship, Principle of relative dating,principle of horizontality,in some cases absolute dating is required to complete geological history of a region.

Correlation required to know actual age of formation and relative age of formation, in which conditions it was formed, orentation of formation to know.

7)If there are 10000 parent atom after 6 half life it will be 10000/2=5000 in first half life

5000/2=2500 in 2nd half life

2500/2z=1250 in 3 rd half life

1250/2=625 in 4th half life

625/2=312.50in 5th half life

312.50/2=156.25 6th half life

Parent atom will be 156.25 after 6th half life,Daughter atom 10,000-156.25=9843.75 atom

8)C-14 half life 5730 years, it is useful to date sample which has sign of life or fossil especially in the range of Tertiary periods. Where there is fossil especially it is of no use archean and proterozoic age.

4